# CMRNext: Camera to LiDAR Matching in the Wild for Localization and Extrinsic Calibration

### v3

@@ I Introduction
Localization is an essential task for any autonomous robot, as it is a precursor for subsequent safety-critical tasks such as path planning and navigation. The accuracy of Global Navigation Satellite Systems is not sufficient for most of these tasks, and therefore sensors such as Light Detection And Ranging (LiDAR) and cameras are employed to estimate the pose of the robot. While LiDAR-based localization methods [1, 2] achieve the best performance in terms of accuracy and robustness, the comparatively high price of LiDAR sensors hinders their widespread adoption. On the other hand, cameras are inexpensive and already widely available in consumer vehicles for Advanced Driver Assistance Systems (ADAS) applications. However, the performance of camera-based localization methods [3, 4, 5, 6] is still not comparable to LiDAR-based localization techniques, especially when the lighting and weather conditions during deployment differ from those recorded while mapping.
@@
Consequently, monocular localization in LiDAR maps has gained considerable attention from both academia and industry in recent years, as it combines the high accuracy and robustness of LiDAR-based localization, while only requiring a camera sensor during deployment. The increasing availability of LiDAR maps from mapping companies such as HERE and TomTom has further contributed to the popularity of this approach. Once a LiDAR map is available, either obtained from a mapping company or generated using recent LiDAR Simultaneouos Localization And Mapping (SLAM) algorithms [7, 2, 8], a vehicle can be localized within this map using a single camera. Considering the low cost of cameras, compared to LiDAR sensors, this approach is particularly appealing for large-scale deployment, where the cost of the final system is a key factor. While traditionally this task has been addressed by first generating a 3D representation from the camera image and then matching it against the LiDAR map [9], the recent advancement of Deep Neural Network (DNN)-based methods for multimodal representation learning [10, 11, 12] has enabled the matching to be performed directly in the feature space.
@@
Most of the existing DNN-based approaches for monocular localization in LiDAR maps, however, follow the end-to-end learning paradigm, where the network directly predicts the metric pose of the camera given an input image and a representation of the LiDAR map. While this paradigm has the advantage of being simple to train and usually achieves good performance when evaluated on images from the same camera used during training, it has the major drawback of being dependent on the specific camera’s intrinsic parameters. Consequently, the network has to be retrained for every new camera, and cannot generalize to different sensor setups. This is even worse when these approaches are applied to the task of extrinsic calibration, where the goal is to estimate the relative pose between the camera and the LiDAR sensor. In this case, end-to-end approaches trained on a single camera-LiDAR pair typically tend to overfit to the specific intrinsic and extrinsic parameters of the dataset and fail to generalize even when evaluated on a different camera setup in the same dataset. For example, a network trained in an end-to-end manner on the left camera of the KITTI dataset [13] and evaluated on the right camera, will still predict the pose of the left camera. This shows that these networks just learn the specific extrinsic parameters of the camera, and thus cannot be used for extrinsic calibration in any real-world application.
@@
In this paper, we propose CMRNext, a novel DNN-based approach for camera-LiDAR matching that is independent of any sensor-specific parameter. Our approach can be employed for monocular localization in LiDAR-maps, as well as camera-LiDAR extrinsic calibration. Moreover, CMRNext is able to generalize to different environments and different sensor setups, including different intrinsic and extrinsic parameters. This is achieved by decoupling the matching step from the metric pose estimation step. In particular, we first train a Convolutional Neural Network (CNN) to predict dense matches between image pixels and 3D points in the LiDAR point clouds, together with their respective uncertainties. These matches are then used to estimate the pose of the camera using a traditional Perspective-n-Points (PnP) algorithm. Hence, the network only reasons at the pixel level, and is thus independent of any metric information. The intrinsic parameters of the camera are instead given as input to the PnP algorithm, and therefore the network can be used with different cameras without needing any retraining. We perform comprehensive experimental evaluations that demonstrate that CMRNext outperforms existing methods and achieves state-of-the-art performance on both monocular localization in LiDAR-maps and extrinsic calibration on three publicly available datasets. Moreover, we evaluate our approach on three in-house robotic platforms: a self-driving perception car, a quadruped robot, and a quadcopter, and show that CMRNext can generalize to very different robotic platforms and sensor setups without any retraining or fine-tuning.
@@ II Related Works
Due to the intrinsically different nature of cameras and LiDAR sensors, standard approaches for monocular localization in LiDAR-maps first convert one modality into the other, such that the matching can be performed in the same space. For example, Caselitz et al. [9] propose to generate a 3D point cloud from the image stream using visual odometry and local bundle adjustment, and then continuously match the reconstructed points with the LiDAR map to track the camera position. subsequent works further improved this approach by using an updatable scale estimator to reduce the scale-drift problem [14] or by extending it to multiple camera setups [15]. Conversely, the match can be performed in the visual space, by projecting the LiDAR map into the image plane and comparing the intensities produced by the LiDAR sensor with a grayscale image using normalized mutual information [16] or normalized information distance [17]. The former family of approaches, however, is based on a point cloud generated from monocular visual odometry, which is noisy and requires a continuous estimation of the scale factor. The latter family of approaches, instead, is based on the assumption that LiDAR intensities and grayscale images are visually similar, which is not always true, especially considering that intensities vary greatly across different LiDAR sensor models and manufacturers [18].
Recently, DNN-based approaches have been proposed to directly match the two modalities in the feature space, taking advantage of recent advancements in robust feature learning. Prior works on DNN-based metric monocular localization [19] implicitly learn the map during the training stage, and thus can only be employed in the same area used for training. Our previous work CMRNet [20] was the first DNN-based method to use the LiDAR map as additional input, such that the network learns to match the camera image with the LiDAR map. Thereby, CMRNet was able to localize a camera image in any environment for which a LiDAR map is available without any retraining or fine-tuning. Based on CMRNet, HyperMap [21] proposes a method to compress the LiDAR map without sacrificing localization performances, while POET [22] replaced the fully-connected layers of CMRNet with a Transformer-based pose regressor. These approaches, however, while being able to generalize to different environments, are still dependent on the specific camera used during training. Concurrently, a parallel branch of research proposed to match keypoints across images and LiDAR maps. 2D3D-MatchNet [23] was the first proposed method in this category, and it leveraged DNN-based feature extractors and metric learning. Few recent works have since been proposed [24, 25], and although they are intrinsically camera and map-agnostic, they either focus on indoor environments or achieve performance that are not suitable for mobile robotics applications when evaluated in outdoor environments.
@@
The extrinsic calibration between LiDAR and camera sensors is typically performed using target-based methods, where an ad-hoc calibration board is placed in the environment, and the relative pose between the two sensors is estimated by detecting the board in both modalities. For example, common target boards used for this task are one or multiple checkerboards on planar boards [28, 29], ordinary boxes [30] or a planar board with circular holes [31]. Although target-based methods are widely used, they require a data recording specifically for this purpose, which should be repeated periodically to account for possible changes in the extrinsic parameters due to vibrations, temperature variations, and manual repositioning of the sensors. Early attempts at targetless extrinsic calibration used visual similarity measures, such as mutual information [32], contours [33], or edges [34], to estimate the relative pose between the sensors. These methods, however, are sensitive to appearance changes, such as lighting and weather conditions, and thus require a careful tuning of the parameters.
@@
In recent years, DNN-based approaches have been proposed to estimate the extrinsic parameters between LiDAR and camera sensors. RegNet [35] was a pioneering method in this field, and it used a CNN that processed a camera image and a LiDAR projection to directly estimate the relative rotation and translation between the two sensors. Many subsequent works follow a similar approach and propose different network architectures to improve the performance [36, 37]. More recently, taking inspiration from CMRNet, LCCNet [38] proposed to leverage the cost volume usually employed in optical flow networks [39] instead of the simple concatenation to aggregate features from the two modalities. These approaches, however, are intrinsically dependent on the specific camera-LiDAR pair used during training, and thus cannot be used for extrinsic calibration in any other sensor setup. In fact, all the aforementioned methods are evaluated only on the same camera-LiDAR pair used during training, which is, in our opinion, an improper evaluation protocol for the extrinsic calibration task.
Following CMRNet++ for monocular localization, a few camera-agnostic approaches based on flow estimation between camera and LiDAR projection have been proposed for extrinsic calibration [40, 41]. Although these approaches are independent of the camera’s intrinsic parameters, when evaluated on a different robotic platform, they still need to be fine-tuned on the new camera-LiDAR pair. While fine-tuning is computationally inexpensive compared to the full training required by camera-dependent approaches, requiring ground truth extrinsic calibration to predict the same extrinsic parameters is not a realistic assumption for real-world applications. In this paper, we additionally extend CMRNext for extrinsic calibration, and we demonstrate that our approach can generalize to very different sensor setups in a zero-shot manner, without any retraining or fine-tuning.
@@ III Technical Approach
In this section, we describe our proposed CMRNext for generalizable camera-LiDAR matching, which can be used for tasks such as monocular localization in LiDAR-maps and extrinsic camera-LiDAR calibration. An overview of the proposed method is depicted in Figure 2, and is composed of two main stages: pixel to 3D point matching, followed by 6-DoF pose regression. In the first step, the CNN only focuses on matching at the pixel level instead of metric basis, which makes the network independent of the intrinsic parameters of the camera. These parameters are instead employed in the second step, where traditional computer vision methods are exploited to estimate the pose of the camera, given the matches from the first step. Consequently, after training, CMRNext can also be used with different cameras and maps from those used while training. We first detail the aforementioned stages, followed by the network architecture, the loss function used to train CMRNext, the generation of the ground truth, and the iterative refinement approach employed to improve the predictions.
@@ III-A Matching Step
Due to the sparse nature of LiDAR point clouds, the projection of the map into the virtual image plane can include points that are occluded by other objects. For example, a point that lays behind a building will still be projected in the LiDAR-image if no closer points are present along the same projection line. Therefore, to deal with occlusions, we employ a z-buffer technique followed by an occlusion estimation filter In particular, we reimplemented the occlusion filter proposed in [43], which is based on the insight that a point is considered occluded if some other points obstruct the accessibility of that point. They define the accessibility of a point as the aperture of the maximum cone that connects the point to the pinhole of the camera without intersecting any other points. More formally, for every point which is projected in a neighborhood of size centered around the projection of , we define the normalized vector between the two points as . Then, we compute the angle between this vector and the normalized vector from to the camera pinhole as . We split the occlusion kernel into four sectors and for each sector, we calculate the maximum value. If the sum of the maximum values of the four sectors is greater than a threshold , then the point is considered as visible, otherwise it is considered as occluded. Since the occlusion filter is performed in the image space, it can be efficiently implemented as a GPU kernel. We refer the reader to [43] for more details. The effect of the occlusion filter is shown in Figure 3. Once the inputs to the network (RGB and LiDAR images) have been obtained, for every 3D point in the LiDAR-image, CMRNext estimates which pixel of the RGB image depicts the same world point.
@@ III-B Localization Step
Once CMRNext has been trained, during inference we have the map, i.e., a set of 3D points whose coordinates are known, altogether with their projection in the initial pose , and a set of matching pixels in the RGB image that is computed using the displacements predicted by the CNN given as
@@
While different algorithms to solve the PnP problem have been proposed, they are usually highly sensitive to outliers, due to the least-squares nature of the optimization. For this reason, we employ the robust RANdom SAmple Consensus (RANSAC) estimator [46] by randomly selecting four correspondences, estimating the camera pose based on these correspondences using the EPnP algorithm [47] and calculating how many points have a reprojection error lower than a threshold (inliers). We repeat this process for iterations and we select the pose with the highest number of inliers. Since the LiDAR images can contain tens of thousands of points, computing the reprojection error for each hypothesis is computationally expensive and thus we implemented a GPU-based computation of the reprojection errors.
@@ III-C Network Architecture
We build our pixel-level matching network based on the RAFT architecture [48], originally proposed for optical flow estimation. RAFT outperformed previous methods by a large margin and demonstrated better generalization ability and many of the most recent state-of-the-art approaches are built upon it [49, 50]. Although RAFT shares some similarities with previous coarse-to-fine architectures, such as the well-known PWCNet [39], it presents some key differences that allow it to achieve better performance at a lower computational cost. In coarse-to-fine architectures, a local correlation volume (where each pixel is compared against neighboring pixels) is used to predict a coarse optical flow (e.g., -th of the original image size). The predicted flow is then upsampled by a factor of two and used to warp the features of the image. This process is repeated until the desired resolution is reached (e.g., -th of the original resolution). In contrast, RAFT computes an all-pairs correlation volume, by comparing all pixels in one feature map with every pixel in the other feature map. Although this correlation volume is more computationally and memory intensive than a local correlation volume, it is only computed once and it does not require any warping operation. The optical flow is initialized to zero and iteratively refined via a recurrent Gated Recurrent Unit (GRU) module.
Our modified version of RAFT for camera to LiDAR matching is depicted in Figure 4 and is composed of four main components: feature encoders, context encoder, multi-scale cost volume, and iterative flow refinement. The image encoder first employs a convolutional layer with downsampling by a factor of 2, followed by six residual blocks and downsampling at every other block. The final image feature map has a shape of . The LiDAR and context encoders share the same architecture as the image encoder, except for an additional Fourier feature mapping module. Recent work [51] has shown that Fourier features help coordinate-based DNNs learn high-frequency functions and improve their performance in a wide variety of tasks. Similar feature mapping modules have been used in seminal works such as the positional encoding employed in Transformers [52] and NERFs [53]. In particular, we define the mapping as follows:

### v4

@@ I Introduction
Localization is an essential task for any autonomous robot, as it is a precursor for subsequent safety-critical tasks such as path planning and navigation. The accuracy of Global Navigation Satellite Systems is not sufficient for most of these tasks, and therefore, sensors such as Light Detection And Ranging (LiDAR) and cameras are employed to estimate the pose of the robot. While LiDAR-based localization methods [1, 2] achieve the best performance in terms of accuracy and robustness, the comparatively high price of LiDAR sensors hinders their widespread adoption. On the other hand, cameras are inexpensive and already widely available in consumer vehicles for Advanced Driver Assistance Systems (ADAS) applications. However, the performance of camera-based localization methods [3, 4, 5, 6] is still not comparable to LiDAR-based localization techniques, especially when the lighting and weather conditions during deployment differ from those recorded while mapping.
@@
Consequently, monocular localization in LiDAR maps has gained considerable attention from both academia and industry in recent years, as it combines the high accuracy and robustness of LiDAR-based localization while only requiring a camera sensor during deployment. The increasing availability of LiDAR maps from mapping companies such as HERE and TomTom has further contributed to the popularity of this approach. Once a LiDAR map is available, either obtained from a mapping company or generated using recent LiDAR Simultaneouos Localization And Mapping (SLAM) algorithms [7, 2, 8], a vehicle can be localized within this map using a single camera. Considering the low cost of cameras, compared to LiDAR sensors, this approach is particularly appealing for large-scale deployment, where the cost of the final system is a key factor. While traditionally this task has been addressed by first generating a 3D representation from the camera image and then matching it against the LiDAR map [9], the recent advancement of Deep Neural Network (DNN)-based methods for multimodal representation learning [10, 11, 12] has enabled the matching to be performed directly in the feature space.
@@
Most of the existing DNN-based approaches for monocular localization in LiDAR maps, however, follow the end-to-end learning paradigm, where the network directly predicts the metric pose of the camera given an input image and a representation of the LiDAR map. While this paradigm has the advantage of being simple to train and usually achieves good performance when evaluated on images from the same camera used during training, it has the major drawback of being dependent on the specific camera’s intrinsic parameters. Consequently, the network has to be retrained for every new camera and cannot generalize to different sensor setups. This is even worse when these approaches are applied to the task of extrinsic calibration, where the goal is to estimate the relative pose between the camera and the LiDAR sensor. In this case, end-to-end approaches trained on a single camera-LiDAR pair typically tend to overfit to the specific intrinsic and extrinsic parameters of the dataset and fail to generalize even when evaluated on a different camera setup in the same dataset. For example, a network trained in an end-to-end manner on the left camera of the KITTI dataset [13] and evaluated on the right camera will still predict the pose of the left camera. This shows that these networks just learn the specific extrinsic parameters of the camera and thus cannot be used for extrinsic calibration in any real-world application.
@@
In this paper, we propose CMRNext, a novel DNN-based approach for camera-LiDAR matching that is independent of any sensor-specific parameter. Our approach can be employed for monocular localization in LiDAR-maps, as well as camera-LiDAR extrinsic calibration. Moreover, CMRNext is able to generalize to different environments and different sensor setups, including different intrinsic and extrinsic parameters. This is achieved by decoupling the matching step from the metric pose estimation step. In particular, we first train a Convolutional Neural Network (CNN) to predict dense matches between image pixels and 3D points in the LiDAR point clouds, together with their respective uncertainties. These matches are then used to estimate the pose of the camera using a traditional Perspective-n-Points (PnP) algorithm. Hence, the network only reasons at the pixel level and is thus independent of any metric information. Although the intrinsic parameters of the camera are still required to be known, they are provided as input to the PnP algorithm, and therefore, the network can be used with different cameras without needing any retraining. We perform comprehensive experimental evaluations that demonstrate that CMRNext outperforms existing methods and achieves state-of-the-art performance on both monocular localization in LiDAR-maps and extrinsic calibration on three publicly available datasets. Moreover, we evaluate our approach on three in-house robotic platforms: a self-driving perception car, a quadruped robot, and a quadcopter, and show that CMRNext can generalize to very different robotic platforms and sensor setups without any retraining or fine-tuning.
@@ II Related Works
Due to the intrinsically different nature of cameras and LiDAR sensors, standard approaches for monocular localization in LiDAR-maps first convert one modality into the other, such that the matching can be performed in the same space. For example, Caselitz et al. [9] propose to generate a 3D point cloud from the image stream using visual odometry and local bundle adjustment, and then continuously match the reconstructed points with the LiDAR map to track the camera position. subsequent works further improved this approach by using an updatable scale estimator to reduce the scale-drift problem [14] or by extending it to multiple camera setups [15]. Conversely, the match can be performed in the visual space, by projecting the LiDAR map into the image plane and comparing the intensities produced by the LiDAR sensor with a grayscale image using normalized mutual information [16, 17] or normalized information distance [18]. The former family of approaches, however, is based on a point cloud generated from monocular visual odometry, which is noisy and requires a continuous estimation of the scale factor. The latter family of approaches, instead, is based on the assumption that LiDAR intensities and grayscale images are visually similar, which is not always true, especially considering that intensities vary greatly across different LiDAR sensor models and manufacturers [19].
Recently, DNN-based approaches have been proposed to directly match the two modalities in the feature space, taking advantage of recent advancements in robust feature learning. Prior works on DNN-based metric monocular localization [20] implicitly learn the map during the training stage and thus can only be employed in the same area used for training. Our previous work CMRNet [21] was the first DNN-based method to use the LiDAR map as additional input, such that the network learns to match the camera image with the LiDAR map. Thereby, CMRNet was able to localize a camera image in any environment for which a LiDAR map is available without any retraining or fine-tuning. Based on CMRNet, HyperMap [22] proposes a method to compress the LiDAR map without sacrificing localization performances, while POET [23] replaced the fully connected layers of CMRNet with a Transformer-based pose regressor. These approaches, however, while being able to generalize to different environments, are still dependent on the specific camera used during training. Concurrently, a parallel branch of research proposed to match keypoints across images and LiDAR maps. 2D3D-MatchNet [24] was the first proposed method in this category, and it leveraged DNN-based feature extractors and metric learning. Few recent works have since been proposed [25, 26], and although they are intrinsically camera and map-agnostic, they either focus on indoor environments or achieve performance that are not suitable for mobile robotics applications when evaluated in outdoor environments.
@@
The extrinsic calibration between LiDAR and camera sensors is typically performed using target-based methods, where an ad-hoc calibration board is placed in the environment, and the relative pose between the two sensors is estimated by detecting the board in both modalities. For example, common target boards used for this task are one or multiple checkerboards on planar boards [29, 30], ordinary boxes [31] or a planar board with circular holes [32]. Although target-based methods are widely used, they require a data recording specifically for this purpose, which should be repeated periodically to account for possible changes in the extrinsic parameters due to vibrations, temperature variations, and manual repositioning of the sensors. Early attempts at targetless extrinsic calibration used visual similarity measures, such as mutual information [33, 17], contours [34], or edges [35], to estimate the relative pose between the sensors. These methods, however, are sensitive to appearance changes, such as lighting and weather conditions, and thus require a careful tuning of the parameters.
@@
In recent years, DNN-based approaches have been proposed to estimate the extrinsic parameters between LiDAR and camera sensors. RegNet [36] was a pioneering method in this field, and it used a CNN that processed a camera image and a LiDAR projection to directly estimate the relative rotation and translation between the two sensors. Many subsequent works follow a similar approach and propose different network architectures to improve the performance [37, 38]. More recently, taking inspiration from CMRNet, LCCNet [39] proposed to leverage the cost volume usually employed in optical flow networks [40] instead of the simple concatenation to aggregate features from the two modalities. These approaches, however, are intrinsically dependent on the specific camera-LiDAR pair used during training and thus cannot be used for extrinsic calibration in any other sensor setup. In fact, all the aforementioned methods are evaluated only on the same camera-LiDAR pair used during training, which is, in our opinion, an improper evaluation protocol for the extrinsic calibration task. Similarly, RGKCNet [41] associates only keypoints rather than the entire point cloud, relying on a dedicated keypoint extraction network. However, this requires a sufficient number of evenly distributed keypoints in the scene to ensure parameter observability. An alternative approach is to align features produced by a deep network, followed by a subsequent optimization step that makes the method generalizable across different sensor setups and scenarios [42]. Following CMRNet++ for monocular localization, a few camera-agnostic approaches based on flow estimation between camera and LiDAR projection have been proposed for extrinsic calibration [43, 44]. Although these approaches are independent of the camera’s intrinsic parameters, when evaluated on a different robotic platform, they still need to be fine-tuned on the new camera-LiDAR pair. While fine-tuning is computationally inexpensive compared to the full training required by camera-dependent approaches, requiring ground truth extrinsic calibration to predict the same extrinsic parameters is not a realistic assumption for real-world applications. In this paper, we additionally extend CMRNext for extrinsic calibration, and we demonstrate that our approach can generalize to very different sensor setups in a zero-shot manner without any retraining or fine-tuning.
@@ III Technical Approach
In this section, we describe our proposed CMRNext for generalizable camera-LiDAR matching, which can be used for tasks such as monocular localization in LiDAR-maps and extrinsic camera-LiDAR calibration. An overview of the proposed method is depicted in Figure 2, and is composed of two main stages: pixel to 3D point matching, followed by 6-DoF pose regression. In the first step, the CNN only focuses on matching at the pixel level instead of metric basis, which makes the network independent of the intrinsic parameters of the camera. These parameters are instead employed in the second step, where traditional computer vision methods are exploited to estimate the pose of the camera, given the matches from the first step. Consequently, after training, CMRNext can also be used with different cameras and maps from those used while training. Although the method is agnostic to the camera’s intrinsic parameters, these parameters still need to be provided as input, and we assume them to be accurate. We first detail the aforementioned stages, followed by the network architecture, the loss function used to train CMRNext, the generation of the ground truth, and the iterative refinement approach employed to improve the predictions.
@@ III-A Matching Step
Due to the sparse nature of LiDAR point clouds, the projection of the map into the virtual image plane can include points that are occluded by other objects. For example, a point that lays behind a building will still be projected in the LiDAR-image if no closer points are present along the same projection line. Therefore, to deal with occlusions, we employ a z-buffer technique followed by an occlusion estimation filter In particular, we reimplemented the occlusion filter proposed in [46], which is based on the insight that a point is considered occluded if some other points obstruct the accessibility of that point. They define the accessibility of a point as the aperture of the maximum cone that connects the point to the pinhole of the camera without intersecting any other points. More formally, for every point which is projected in a neighborhood of size centered around the projection of , we define the normalized vector between the two points as . Then, we compute the angle between this vector and the normalized vector from to the camera pinhole as . We split the occlusion kernel into four sectors, and for each sector, we calculate the maximum value. If the sum of the maximum values of the four sectors is greater than a threshold , then the point is considered as visible, otherwise it is considered as occluded. Since the occlusion filter is performed in the image space, it can be efficiently implemented as a GPU kernel. We refer the reader to [46] for more details. The effect of the occlusion filter is shown in Figure 3. Once the inputs to the network (RGB and LiDAR images) have been obtained, for every 3D point in the LiDAR-image, CMRNext estimates which pixel of the RGB image depicts the same world point.
@@ III-B Localization Step
Once CMRNext has been trained, during inference, we have the map, i.e., a set of 3D points whose coordinates are known, altogether with their projection in the initial pose , and a set of matching pixels in the RGB image that is computed using the displacements predicted by the CNN given as
Although CMRNext predicts dense correspondences, even for pixels without an associated 3D LiDAR point, we first filter out invalid pixels by using the defined in Equation 6
@@
While different algorithms to solve the PnP problem have been proposed, they are usually highly sensitive to outliers due to the least-squares nature of the optimization. For this reason, we employ the robust RANdom SAmple Consensus (RANSAC) estimator [49] by randomly selecting four correspondences, estimating the camera pose based on these correspondences using the EPnP algorithm [50] and calculating how many points have a reprojection error lower than a threshold (inliers). We repeat this process for iterations and we select the pose with the highest number of inliers. Since the LiDAR images can contain tens of thousands of points, computing the reprojection error for each hypothesis is computationally expensive, and thus, we implemented a GPU-based computation of the reprojection errors.
@@ III-C Network Architecture
We build our pixel-level matching network based on the RAFT architecture [51], originally proposed for optical flow estimation. RAFT outperformed previous methods by a large margin and demonstrated better generalization ability, and many of the most recent state-of-the-art approaches are built upon it [52, 53]. Although RAFT shares some similarities with previous coarse-to-fine architectures, such as the well-known PWCNet [40], it presents some key differences that allow it to achieve better performance at a lower computational cost. In coarse-to-fine architectures, a local correlation volume (where each pixel is compared against neighboring pixels) is used to predict a coarse optical flow (e.g., -th of the original image size). The predicted flow is then upsampled by a factor of two and used to warp the features of the image. This process is repeated until the desired resolution is reached (e.g., -th of the original resolution). In contrast, RAFT computes an all-pairs correlation volume by comparing all pixels in one feature map with every pixel in the other feature map. Although this correlation volume is more computationally and memory intensive than a local correlation volume, it is only computed once, and it does not require any warping operation. The optical flow is initialized to zero and iteratively refined via a recurrent Gated Recurrent Unit (GRU) module.
Our modified version of RAFT for camera to LiDAR matching is depicted in Figure 4 and differs from the original architecture by using separate feature encoders for camera and LiDAR inputs, using the LiDAR image in the context encoder with added positional encoding, integrating uncertainty estimation in the update block, and replacing the L1 loss with a masked Negative Log-Likelihood loss. We will analyze all components in the following. The image encoder first employs a convolutional layer with downsampling by a factor of 2, followed by six residual blocks and downsampling at every other block. The final image feature map has a shape of . The LiDAR and context encoders share the same architecture as the image encoder, except for an additional Fourier feature mapping module. Recent work [54] has shown that Fourier features help coordinate-based DNNs learn high-frequency functions and improve their performance in a wide variety of tasks. Similar feature mapping modules have been used in seminal works such as the positional encoding employed in Transformers [55] and NERFs [56]. In particular, we define the mapping as follows: